The supply of wheat is given by the following equation: Qws = -6 + 4Pw - 2Pc - P, where Qws is the quantity o wheat supplied, in millions of bushes; Pw is the price of wheat per bushel; Pc is the price of corn per bushel; and Pf is the price of tractor fuel per gallon. When corn sells for $4 a bushel and fuel sells for 42 a gallon: What is the equation for the supply curve?uel per gallon
Answer #1

Qw = -6 +4Pw -2Pc -Pf

Pc= $4

Pf = $42

Qw = -6 +4Pw -2(4) - (42)

= -6 + 4Pw -8 -42

Qw =-56 +4 Pw

Equation of the supply curve : Qw = -56 +4 Pw